OpenStudy (anonymous):

the sales tax on a metal is $6.00 meal is $0.36. What percent sales tax is being charged?

OpenStudy (aum):

\(\text{Percentage Sales Tax} = \Large \frac{\text{Sales Tax}}{\text{Cost of item}} \times 100 \text{%} = \frac{0.36}{6.00} \times 100 = ?~\text{%}\)

OpenStudy (triciaal):

another way of saying what @aum did percent is per 100 let the rate = x set up a ratio x/100= 0.36/6.00 solve for x